# On polynomials associated to Voronoi diagrams of point sets and crossing numbers

### v2

@@ 1 Introduction
The rectilinear crossing number of a point set $S$, $\overline{cr}(S)$, is the number of pairwise edge crossings of the complete graph $K_{n}$ when drawn with straight-line segments on $S$, i.e. the vertices of $K_{n}$ are the points of $S$. Equivalently, $\overline{cr}(S)$ is the number of convex quadrilaterals with vertices in $S$. We denote $\overline{cr}(S)$ as $\alpha{{n}\choose{4}}$, with $0\leq\alpha\leq 1$. Note that for $S$ in convex position, $\alpha=1$. The rectilinear crossing number problem consists in, for each $n$, finding the minimum value of $\overline{cr}(S)$ among all sets $S$ of $n$ points, no three of them collinear. This minimum is commonly denoted as $\overline{cr}(K_{n})$. The limit of ${\overline{cr}(K_{n})}/{{n}\choose{4}}$, when $n$ tends towards infinity, is the so-called rectilinear crossing number constant $\alpha^{*}$. This problem is solved for $n\leq 27$ and $n=30$, and the current best bound for the rectilinear crossing number constant is $\alpha^{*}>0,37997$, for more information, see the survey of Ábrego et al. (2013) and the web page Aichholzer (a). A fruitful approach to the rectilinear crossing number problem is proving bounds on the numbers of $j$-edges and of $(\leq k)$-edges of $S$Ábrego et al. (2012); Ábrego and Fernández-Merchant (2017); Aichholzer et al. (2007); Balogh and Salazar (2006); Lovász et al. (2004). An (oriented) $j$-edge of $S$ is a directed straight line $\ell$ passing through two points of $S$ such that the open half-plane bounded by $\ell$ and on the right of $\ell$ contains exactly $j$ points of $S$. The number of $j$-edges of $S$ is denoted by $e_{j}$, and $E_{\leq k}=\sum_{j=0}^{k}e_{j}$ is the number of $(\leq k)$-edges. We then consider the $E_{\leq k}$ polynomial $p_{E}(z)=\sum_{k=0}^{n-3}E_{\leq k}z^{k}$, which also encodes information on higher order Voronoi diagrams, since the number of $j$-edges $e_{j}$ is the number of unbounded cells of the order-$(j+1)$ Voronoi diagram of $S$, see e.g. Proposition 30 in Claverol et al. (2021a). Note that $p_{E}(z)$ has no term $E_{\leq n-2}$.
@@
Figure 1: Left: 1591-th entry of the order type database for 8 points, from  Aichholzer (b). With complex stream plots of its Voronoi polynomial (center): $p_{V}(z)=10+23z+27z^{2}+24z^{3}+17z^{4}+9z^{5}+2z^{6}$, and its $E_{\leq k}$ polynomial (right): $p_{E}(z)=4+13z+22z^{2}+34z^{3}+43z^{4}+52z^{5}$; roots are red points.
For a point set $S$, we show that $\overline{cr}(S)$ appears in the first derivatives of these three polynomials when evaluated at $z=1$ and, in addition, we obtain appealing formulas for $\overline{cr}(S)$ in terms of the roots of the polynomials. Motivated by this, we study the location of such roots, showing several bounds on their modulus. As a particular result, we also prove that the roots of the Voronoi polynomial lie on the unit circle, $\{z:|z|=1\}$, if and only if $S$ is in convex position. Polynomials with zeros on the unit circle have been studied for instance in Chapoton and Han (2020); Chen (1995); Lakatos and Losonczi (2004); Lalín and Smyth (2013).
@@ 2 Known relations
  • For any point set $S$, and $30\leq k\leq n-$ 3, it holds that, see Ardila (2004); Clarkson and Shor (1989); Claverol et al. (2021a); Lee (1982); Lindenbergh (2003),
@@
    This was essentially also obtained in Urrutia (February 2004), though not stated in terms of $\overline{cr}(S)$.
@@
  • For $k\leq\frac{n-3}{2}$ it holds that, see Lemma 3.1 in Claverol et al. (2021b),
@@
  • For every set $S$ of $n$ points in general position, the relation between $E_{\leq k}$ and $c_{k}$ is, see e.g. Property 33 in Claverol et al. (2021a),
@@
    For a point set $S$ in convex position, we have equality in Equations (5) and (6) which can be derived from Equations (7) and $E_{\leq k}=(k+1)n$ for $S$ in convex position, and therefore,
@@
    Then, the number of vertices of $V_{k}(S)$, for $S$ in convex position fulfills, see e.g. Property 34, Equation (4) in Claverol et al. (2021a),
@@ 3 Properties of the Voronoi, circle and $E_{\leq_{k}}$ polynomials
For every set $S$ of $n$ points in general position, the Voronoi polynomial $p_{V}(z)=\sum_{k=1}^{n-1}v_{k}z^{k-1}$ satisfies:
@@ 4 On the roots of Voronoi, circle and $E_{\leq_{k}}$ polynomials
Figure 3: Left: Best known example for minimizing $\overline{cr}(S)$ for $|S|=100$, from  Aichholzer (a). Roots of its Voronoi (center), and $E_{\leq k}$ (right) polynomials, with circles illustrating, respectively, the bounds of Theorems 13 and 18.

### v3

@@ 1 Introduction
The rectilinear crossing number of a point set $S$, $\overline{cr}(S)$, is the number of pairwise edge crossings of the complete graph $K_{n}$ when drawn with straight-line segments on $S$, i.e. the vertices of $K_{n}$ are the points of $S$. Equivalently, $\overline{cr}(S)$ is the number of convex quadrilaterals with vertices in $S$. We denote $\overline{cr}(S)$ as $\alpha{{n}\choose{4}}$, with $0\leq\alpha\leq 1$. Note that for $S$ in convex position, $\alpha=1$. The rectilinear crossing number problem consists in, for each $n$, finding the minimum value of $\overline{cr}(S)$ among all sets $S$ of $n$ points, no three of them collinear. This minimum is commonly denoted as $\overline{cr}(K_{n})$. The limit of ${\overline{cr}(K_{n})}/{{n}\choose{4}}$, when $n$ tends towards infinity, is the so-called rectilinear crossing number constant $\alpha^{*}$. This problem is solved for $n\leq 27$ and $n=30$, and the current best bound for the rectilinear crossing number constant is $\alpha^{*}>0,37997$; for more information, see the survey of Ábrego et al. (2013) and the web page Aichholzer (2006a). A fruitful approach to the rectilinear crossing number problem is proving bounds on the numbers of $j$-edges and of $(\leq k)$-edges of $S$, see Ábrego et al. (2012); Ábrego and Fernández-Merchant (2017); Aichholzer et al. (2007); Balogh and Salazar (2006); Lovász et al. (2004). An (oriented) $j$-edge of $S$ is a directed straight line $\ell$ passing through two points of $S$ such that the open half-plane bounded by $\ell$ and on the right of $\ell$ contains exactly $j$ points of $S$. The number of $j$-edges of $S$ is denoted by $e_{j}$, and $E_{\leq k}=\sum_{j=0}^{k}e_{j}$ is the number of $(\leq k)$-edges. We then consider the $E_{\leq k}$ polynomial $p_{E}(z)=\sum_{k=0}^{n-3}E_{\leq k}z^{k}$, which also encodes information on higher order Voronoi diagrams, since the number of $j$-edges $e_{j}$ is the number of unbounded cells of the order-$(j+1)$ Voronoi diagram of $S$, see e.g. Proposition 30 in Claverol et al. (2024). Note that $p_{E}(z)$ has no term $E_{\leq n-2}$.
@@
Figure 1: Left: 1591-th entry of the order type database for 8 points, from  Aichholzer (2006b). With complex stream plots of its Voronoi polynomial (center): $p_{V}(z)=10+23z+27z^{2}+24z^{3}+17z^{4}+9z^{5}+2z^{6}$, and its $E_{\leq k}$ polynomial (right): $p_{E}(z)=4+13z+22z^{2}+34z^{3}+43z^{4}+52z^{5}$; roots are red points.
For a point set $S$, we show that $\overline{cr}(S)$ appears in the first derivatives of these three polynomials when evaluated at $z=1$ and, in addition, we obtain appealing formulas for $\overline{cr}(S)$ in terms of the roots of the polynomials. Motivated by this, we study the location of such roots, showing several bounds on their modulus. As a particular result, we also prove that the roots of the Voronoi polynomial lie on the unit circle, $\{z:|z|=1\}$, if, and only if, $S$ is in convex position. Polynomials with zeros on the unit circle have been studied for instance in Chapoton and Han (2020); Chen (1995); Lakatos and Losonczi (2004); Lalín and Smyth (2013).
@@ 2 Known relations
  • For any point set $S$, and $30\leq k\leq n-$ 3, it holds that, see Ardila (2004); Clarkson and Shor (1989); Claverol et al. (2024); Lee (1982); Lindenbergh (2003),
@@
    This was essentially also obtained in Urrutia (2004), though not stated in terms of $\overline{cr}(S)$.
@@
  • For $k\leq\frac{n-3}{2}$ it holds that, see Lemma 3.1 in Claverol et al. (2021),
@@
  • For every set $S$ of $n$ points in general position, the relation between $E_{\leq k}$ and $c_{k}$ is, see e.g.
    Property 33 in Claverol et al. (2024),
@@
    For a point set $S$ in convex position, we have equality in Equations (5) and (6) which can be derived from Equations (7) and $E_{\leq k}=(k+1)n$ for $S$ in convex position, and therefore
@@
    Then, the number of vertices of $V_{k}(S)$, for $S$ in convex position fulfills, see e.g. Property 34, Equation (4) in Claverol et al. (2024),
@@ 3 Properties of the Voronoi, circle and $E_{\leq_{k}}$ polynomials
For every set $S$ of $n$ points in general position, the Voronoi polynomial
$p_{V}(z)=\sum_{k=1}^{n-1}v_{k}z^{k-1}$ satisfies:
@@ 4 On the roots of Voronoi, circle and $E_{\leq_{k}}$ polynomials
Figure 3: Left: Best known example for minimizing $\overline{cr}(S)$ for $|S|=100$, from  Aichholzer (2006a). Roots of its Voronoi (center), and $E_{\leq k}$ (right) polynomials, with circles illustrating, respectively, the bounds of Theorems 13 and 18.